The table below shows the weight of 21 students (on the left) and the corresponding number of students with that weight (on the right). Compute the Mean of the data below.



Weight No. of Students


34 5


39 8


43 2


45 6


1
Expert's answer
2022-02-17T13:16:35-0500

Represent the weight by "x" and the number of students by "f".

Now,

"x" "f" "fx"

34 5 170


39 8 312


43 2 86


45 6 270

The mean, "\bar x={\sum fx\over \sum f}={838\over21}=39.90(2dp)"